# FineReason: Evaluating and Improving LLMs’ Deliberate Reasoning through Reflective Puzzle Solving

### v1

@@ A.4 Additional Experimental Results
Current state: [[4, 1, 6, 9, 7, 2, 8, 3, 5], [7, 2, 3, 1, 8, 5, 6, 9, 4], [5, 9, 8, 3, 4, 6, 2, 1, 7], [6, 3, 5, 4, 1, 9, 7, 2, 8], [1, 8, 9, 2, 6, 7, 4, 5, 3], [2, 4, 7, 5, 3, 8, 1, 6, 9], [8, 7, 2, 6, 9, 3, 5, 4, 1], [3, 6, 0, 0, 5, 0, 0, 7, 2], [0, 0, 0, 0, 0, 4, 3, 8, 0]]
Explored next state that leads to an unsolvable path: [[4, 1, 6, 9, 7, 2, 8, 3, 5], [7, 2, 3, 1, 8, 5, 6, 9, 4], [5, 9, 8, 3, 4, 6, 2, 1, 7], [6, 3, 5, 4, 1, 9, 7, 2, 8], [1, 8, 9, 2, 6, 7, 4, 5, 3], [2, 4, 7, 5, 3, 8, 1, 6, 9], [8, 7, 2, 6, 9, 3, 5, 4, 1], [3, 6, 1, 0, 5, 0, 0, 7, 2], [0, 0, 0, 0, 0, 4, 3, 8, 0]]
@@
State Transition
You are given an initial Sudoku puzzle S(0), followed by a sequence of progressive states leading to the current state S(i). Alongside each state, its solvability status L(*) is given. Your task is to determine the next state by making exactly one move, ensuring progress toward a valid solution. A valid Sudoku solution requires that each row, column, and 3x3 subgrid contains the numbers 1 to 9 without repetition.

### v2

@@ A.9 Benefits of mixed-training across different reasoning tasks
Current state:
[[4, 1, 6, 9, 7, 2, 8, 3, 5], [7, 2, 3, 1, 8, 5, 6, 9, 4], [5, 9, 8, 3, 4, 6, 2, 1, 7], [6, 3, 5, 4, 1, 9, 7, 2, 8], [1, 8, 9, 2, 6, 7, 4, 5, 3], [2, 4, 7, 5, 3, 8, 1, 6, 9], [8, 7, 2, 6, 9, 3, 5, 4, 1], [3, 6, 0, 0, 5, 0, 0, 7, 2], [0, 0, 0, 0, 0, 4, 3, 8, 0]]
Explored next state that leads to an unsolvable path:
[[4, 1, 6, 9, 7, 2, 8, 3, 5], [7, 2, 3, 1, 8, 5, 6, 9, 4], [5, 9, 8, 3, 4, 6, 2, 1, 7], [6, 3, 5, 4, 1, 9, 7, 2, 8], [1, 8, 9, 2, 6, 7, 4, 5, 3], [2, 4, 7, 5, 3, 8, 1, 6, 9], [8, 7, 2, 6, 9, 3, 5, 4, 1], [3, 6, 1, 0, 5, 0, 0, 7, 2], [0, 0, 0, 0, 0, 4, 3, 8, 0]]
@@
Sudoku State Transition
You are given an initial Sudoku puzzle S(0), followed by a sequence of progressive states leading to the current state S(i). Alongside each state, its solvability status L(*) is given.
Your task is to determine the next state by making exactly one move, ensuring progress toward a valid solution. A valid Sudoku solution requires that each row, column, and 3x3 subgrid contains the numbers 1 to 9 without repetition.